OpenStudy (anonymous):

whats 13 1/9 divided by 5 9/10

OpenStudy (anonymous):

\[\frac{\frac{(13*9)+1}{9}}{\frac{(5*10)+9}{10}}=\frac{118}{9}*\frac{10}{59}=\frac{20}{9}\]\]

OpenStudy (anonymous):

118/9 ÷ 59/10 = 118/9 × 10/59 =20/9